OpenStudy (anonymous):

Explain the steps that are necessary to solve 4x - 2(x + 3) = x + 1

OpenStudy (anonymous):

Explain the steps that are necessary to solve 4x - 2(x + 3) = x + 1 Can you explain using complete sentences? i really need your help with this problem thanks. -Z.J

OpenStudy (anonymous):

you want to get all x on one side

OpenStudy (anonymous):

can you solve it for me ? so i can know how to do it?

OpenStudy (anonymous):

First : Get - 2(x + 3) What is the answer of this ?

OpenStudy (petiteme):

4x - 2(x + 3) = x + 1 First, distribute 2 inside the parenthesis so 4x - 2(x) - (2)(3) = x + 1 then perform the operaton 4x - 2x - 6 = x+1 2x-6=x+1 To get x, transpose x to the other side and also -6 so 2x-x = 1+6 x=7

OpenStudy (anonymous):

Solve for x: 4 x-2 (x+3) = x+1 Distribute -2 over x+3. -2 (x+3) = -6-2 x: 4 x+-6-2 x = x+1 Group like terms in 4 x-2 x-6. Grouping like terms, 4 x-2 x-6 = (4 x-2 x)-6: (4 x-2 x)-6 = x+1 Combine like terms in 4 x-2 x. 4 x-2 x = 2 x: 2 x-6 = x+1 Factor the polynomial, 2 x-6. Factor 2 from the polynomial 2 x-6: 2 (x-3) = x+1 Move everything to the left hand side. Subtract x+1 from both sides of 2 (x-3) = x+1: 2 (x-3)-x+1 = (x+1)-x+1 Look for two terms that sum to zero. (x+1)-(x+1) = 0: 2 (x-3)-(x+1) = 0 Distribute 2 over x-3. 2 (x-3) = 2 x-6: 2 x-6-(x+1) = 0 Distribute -1 over x+1. -(x+1) = -1-x: 2 x+-1-x-6 = 0 Group like terms in 2 x-x-6-1. Grouping like terms, 2 x-x-6-1 = (2 x-x)+(-6-1): (2 x-x)+(-6-1) = 0 Combine like terms in 2 x-x. 2 x-x = x: x+(-6-1) = 0 Evaluate -6-1. -6-1 = -7: x+-7 = 0 Isolate terms with x to the left hand side. Add 7 to both sides: x+(7-7) = 7 Look for two terms that sum to zero. 7-7 = 0: Answer: | | x = 7

OpenStudy (anonymous):

Hopefully I helped.

OpenStudy (anonymous):

YES THANK YOU ALL ! :)

OpenStudy (petiteme):

you're welcome :) anytime :)